CYCLES, UNEPLOYMENT, and INFLATION WORKSHEET Student Full Name llica Sooz Days/Time Class Meets rate riven the fellowing data for a country (numbers in millions) Institutionalized 50 Actively-Seeking Work-10 Part-Time Workers- 80 Full-Time Workers-110 Retired Workers- 100 Under 16 years of Age-40 Discouraged Workers- 60 Calculation: Show All Work! 2. Caleulate the GDP Gap given the following data Use Okuns Law) Actual Unemployment Rate = 9,5 % Natural Rate of Unemployment 5 % 3. List the 4 phases of the business cycle in their exact sequence 1) Recession 2) 3) 4) 4. Which is the worst type of unemployment (CIRCLE ONE: Cyclical, Frictional, Structural) WHY? of a) Durable Good- b) Nondurable Good- c) Capital Good- d) Service- e) Which of the above productís) are affected the most by a business cyele? WHY? 6. Which demographic group (CIRCLE ONE: Professional, Women, African American, Ages 16- 19, College Graduate) typically has the highest unemployment rate? WHY?
Answer #1

1) unemployment rate= unemployed*100/labor force

labor force= actively seeking for work + part time and full time workers

= 10+80+110= 200

Unemployment rate= 10*100/200= 5%

2) okun's law-

Output gap= a(unemployment gap)

Unemployment gap= actual unemployment - natural unemployment

= 9.5 - 5= 4.5%

a is assumed to be 2

Output gap= 2(4.5)= 9%